OpenStudy (anonymous):

Measure and record the diameter and height of the cylindrical object you have chosen. Be sure to use the same measurement for both (either centimeters or inches). Apply the formula of a right circular cylinder (V = r2h) to find the volume of the object. (Note: Be sure to find the radius from the diameter measurement by dividing by 2.) Now suppose you knew the volume of this object and the relation of the height to the radius, but did not know the radius. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the radius. Rewrite the form

OpenStudy (anonymous):

r=V/(2*h(r)) --- Where h(r) is h as a function of r.

OpenStudy (anonymous):

Then you would solve for r again after you know h as a function of r.

OpenStudy (anonymous):

For example, lets say the height is always 2r, then r would be... r=V/(2*(2r)) Solving for r, r=sqrt(V/4)